OpenStudy (anonymous):

kyla can cut and split a cord of wood in 5.5 hr. Ronson can cut and split a cord of wood in 6.5 hr. How long will it take them, working together, to cut and split a cord of wood?

OpenStudy (anonymous):

x/5.5 + x/6.5 = 1 12x = 5.5 * 6.5 x = 2.979 hours